Simplify 4(8-6)^2 + 4x3-2x8/4
Answer by benazir.sj@gmail.com(109)

4(8-6)^2 + 4*3-2*8/4
=
4(2)^2+12-2*2
= 4*4+12-4
=16+12-4
=28-4
=24